Last year Lakesha’s Lounge Furniture Corporation had an ROE of 15.6 percent and a dividend payout ratio of 30 percent. What is the sustainable growth rate? What is the sustainable growth rate? (Do not round intermediate calculations. Round your answer to 2 decimal places.)
Answer #1

Retention ratio=1-payout ratio

=(1-0.3)=0.7

Sustainable growth rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]

=(0.156*0.7)/[1-(0.156*0.7)]

=12.26%(Approx).